OpenStudy (anonymous):

Last one i sware Lol. Okay: -6x-3y=10 x=-4+2y

OpenStudy (anonymous):

solve using subsitution

OpenStudy (saifoo.khan):

u want the answers in decimals or in fractions?

OpenStudy (anonymous):

fractions please

OpenStudy (saifoo.khan):

x = -32/15 y = 14/15

OpenStudy (anonymous):

life savers ty =)

OpenStudy (saifoo.khan):

uw. :)